OpenStudy (anonymous):

What is the simplified form of (cube root of)16

OpenStudy (anonymous):

\[2\sqrt[3]{2}\]

OpenStudy (anonymous):

\[2\sqrt[3]{2}\]

OpenStudy (saifoo.khan):

rightt.

OpenStudy (anonymous):

How did you get that though?

OpenStudy (anonymous):

\[\sqrt[3]{2*2*2*2}\]

OpenStudy (saifoo.khan):

CALC.

OpenStudy (anonymous):

I need the steps to write it out.

OpenStudy (anonymous):

\[\sqrt[3]{2*2*2*2}\] \[2\sqrt[3]{2}\]

OpenStudy (anonymous):

these are two steps and i guess they are enough

OpenStudy (anonymous):

Thank you! I think I get it....